OpenStudy (aravindg):

More doubts:

OpenStudy (aravindg):

1.A body with Volume V neither sinks nor floats in a liquid. The vessel containing it falls with an acceleration g/2. Then the volume of solid inside liquid in falling condition is ....

OpenStudy (aravindg):

The answer is V itself. I do not know how that is same.

OpenStudy (vincent-lyon.fr):

Whatever the acceleration of the vessel, if weight is balanced by upthrust at rest, it will always be so. (this is how I understand the wording of your question)

OpenStudy (aravindg):

Then what about the pseudo force acting?

OpenStudy (vincent-lyon.fr):

The pseudo force act equally on the body and on the sustaining water. Since upthrust is equal to "weight" of displaced water, it varies also accordingly.

OpenStudy (vincent-lyon.fr):

*acts

OpenStudy (aravindg):

okay thanks :) 2.Three rods of identical cross sectional area and made from same metal form the sides of an isoceles right angled triangle. ABC rt angled at B. The temperature at points A and B are T and (root2)T in steady state. temperature at C is Tc Find Tc/T

OpenStudy (aravindg):

Here I considered this as a resistance circuit question. Current through AB=Current through BC. But I am not getting the answer in options.

OpenStudy (vincent-lyon.fr):

Tc/T = 2\(\sqrt 2-1\) ?

OpenStudy (aravindg):

Yeah that;s what I got. The answer given is \[\dfrac{\sqrt{3}}{\sqrt{2}+1}\]

OpenStudy (aravindg):

The solution is explained as Heat flowing in BC=Heat flowing in CA

OpenStudy (aravindg):

But for me it looks as they are parallel and hence same current cannot flow.

OpenStudy (vincent-lyon.fr):

Is the triangle open at point B?

OpenStudy (aravindg):

No figure is given.

OpenStudy (aravindg):

Could it be a wrong solution?

OpenStudy (aravindg):

ok

OpenStudy (vincent-lyon.fr):

Typo in your answer? I find \(\dfrac{3}{\sqrt{2}+1}\)

OpenStudy (aravindg):

Nope that was answer given. I can see that if we equate according to their solution there is no chance of getting root(3).

OpenStudy (aravindg):

But how can they equate those currents?

OpenStudy (vincent-lyon.fr):

The thing is Ta and Tb are impose by an external source. So you cannot equate heat flows in AB and BC, only in Bc and CA.

OpenStudy (aravindg):

Sorry I dont get it. The diag in my mind is this: